OpenStudy (anonymous):

how would i solve this? w^2=9

OpenStudy (loser66):

what do you think?

OpenStudy (anonymous):

would i need to get it to this w^2-9=0?

OpenStudy (loser66):

good, then?

OpenStudy (anonymous):

that's where i get confused. would it be (w-1)(w+9)

OpenStudy (loser66):

nope, apply \[a^2 -b^2= (a+b)(a-b)\]your a is w and your b is 3

OpenStudy (anonymous):

so its (w+3)(w-3) that makes sense.

OpenStudy (loser66):

yup

OpenStudy (loser66):

then?

OpenStudy (anonymous):

w=-3 and w=3

OpenStudy (loser66):

medal for you

OpenStudy (anonymous):

thank you!!!

OpenStudy (loser66):

yw